# Learning parameter dependence for Fourier-based option pricing with tensor networks

### v2

@@ I Introduction
In light of this, applications of quantum-inspired algorithms to option pricing have been also studied, as solutions in the present or the near future [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9] learning to the Fourier transform (FT)-based option pricing method [10, 11], following the original proposal in [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [12, 13] as reviews. and recently, it has also been utilized in various fields such as machine learning [14, 15], quantum field theory [16, 17], and differential equations [7, 18]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [19, 20, 17, 21] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
To address these issues, we focus on another usage of tensor networks that can embed parameter dependence [22, 23, 24] to make FT-based option pricing more efficient. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We then use these tensor networks to perform fast option pricing in response to various parameter changes. To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice. When the computational complexity is smaller, our numerical results indicate that our method has a further advantage in terms of computation time, since our method involves only elementary arithmetic contrary to the Monte Carlo method involving more complicated operations. We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
@@ III.1 Fourier transform-based option pricing
where $\Sigma:=(\sigma_{i}\sigma_{j}\rho_{ij}T)_{ij}$ is the covariance matrix of $\vec{x}$ and $\vec{\mu}:=\vec{x}_{0}+\left(rT-\frac{1}{2}\sigma_{1}^{2}T,\cdots,rT-\frac{1}{2}\sigma_{d}^{2}T\right)$. Note that, in Eq. (6), we denote the option price by $V(\vec{p})$, indicating its dependence on the parameter $\vec{p}$ such as the volatilities $\vec{\sigma}=(\sigma_{1},\cdots,\sigma_{d})$ and the present asset prices $\vec{S}_{0}$.
@@
$\phi(\vec{z})=\exp\left(i\sum_{j=1}^{d}z_{j}\mu_{j}-\frac{T}{2}\sum_{j=1}^{d}\sum_{k=1}^{d}\sigma_{j}\sigma_{k}z_{j}z_{k}\rho_{jk}\right).$ (10)
@@
$\hat{v}_{\min}(\vec{z})=-\frac{K^{1+i\sum_{j=1}^{d}z_{j}}}{(-1)^{d}\left(1+i\sum_{j=1}^{d}z_{j}\right)\prod_{j=1}^{d}(iz_{j})}.$ (12)
@@
where $\eta_{j}$ is the integration step size in the $j$-th direction. $\eta_{j}$ is a hyperparameter that must be appropriately determined according to the parameter $\vec{p}$ and thus written as $\eta_{j}(\vec{p})$ in (14). In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec. V for the detail. $\Delta_{\rm vol}:=\Pi_{j=1}^{d}\eta_{j}$ is the volume element.
@@ III.2 Fourier transform-based option pricing with tensor trains
$V(\vec{\sigma},\vec{S}_{0})\simeq\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\tilde{\phi}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}}.$ (15)
Thanks to TCI, we can obtain the approximate TTs avoiding the evaluations of $\phi$ and $\hat{v}$ at all the grid points. Besides, given the TTs, we can compute the sum in (15) as the contraction of two TTs without exponentially many iterations: with the bond dimensions at most $\chi$, the number of multiplications and additions is of order $O(d\chi^{3})$.
@@ III.3 Monte Carlo-based option pricing
IV Learning parameter dependence with tensor trains
@@ IV.1 Outline
Option prices depend on input parameters such as the volatilities $\vec{\sigma}$ and the present asset prices $\vec{S}_{0}$. In the rapidly changing financial market, these input parameters vary from time to time, which causes the change of the option price. Therefore, if we have a function that efficiently outputs an accurate approximation of the option price for various values of the input parameters, it provides a large benefit to practical business.
Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set. Using TCI, we obtain the TTs to approximate $\phi$ and $\hat{v}$, incorporating the parameter dependence of these functions. The outline is illustrated in Fig. 1. Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$. Then, as illustrated in Fig. 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$. We reduce the bond dimensions of these TTs using SVD. Subsequently, we contract adjacent core tensors pairwise and obtain matrix product operators (MPOs). Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$. Having these MPOs, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig. 1 (b). By fixing the index $\vec{k}$ of $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ to the value corresponding to the specified $\vec{p}$, we get the $d$-dimensional TTs $\tilde{\phi}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$. Contracting these TTs as Eq. (15), we get the option price for the specified $\vec{p}$.
@@
Figure 1: Fast option pricing based on TTs proposed in this study. In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD. Then, we contract the tensors associated with $p_{i}$ and $z_{i}$, resulting in MPOs. In (b), we use these MPOs to perform fast option pricing for a specified parameter $\vec{p}$.
@@ IV.2 Computational complexity
The computational complexity of TT-based option pricing, involving computation of the inner product between two TTs, is $O(d\chi^{3}N)$. In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations. Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec. V.
Here, we ignore the computational complexity of learning the TTs by TCI, SVD, and contraction of tensors associated with $p_{i}$ and $z_{i}$ and consider that after we get the MPOs only. This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing. As discussed in Sec. VI, we can reasonably find such a situation in practice.
@@ V.1 Details
Ranges of ${\sigma_{i}}$ and ${S^{i}_{0}}$
With respect to $\vec{p}$, on which the TTs learn the dependence of the functions $\phi$ and $\hat{v}$, we take the two test cases: $\vec{p}$ is the volatilities $\vec{\sigma}$ or the present asset prices $\vec{S}_{0}$. In the proposed method, we need to set the range in the space of $\vec{p}$ and the grid points in it. For each volatility $\sigma_{i}$, we set the range to $\sigma_{i}\in[0.15,0.25]$, where the center $\sigma_{i}=0.2$ is a typical value of Nikkei Stock Average Volatility Index [28] and the range width ±$\pm 0.05$ covers the changes in this index on most days. For each $S_{i,0}$, we set the range to $S_{i,0}\in[90,120]$, which corresponds to the 20% variation of the asset price centered at 100. The lower bound is set to not 80 but 90 because the price of the option we take as an example is negligibly small for $S_{0}<90$. For both $\sigma_{i}$ and $S_{i,0}$, we set 100 equally spaced grid points in the range, and so the total number of the grid points in the space of $\vec{p}$ is $100^{d}$.
@@
The spread of the characteristic function changes depending on the values of the parameters $\vec{p}$ and the number of assets. Accordingly, it is necessary to adjust the grid interval $\eta_{j}$ for integration in the wavenumber space.
For example, consider the case where $\vec{p}=\vec{\sigma}$. From Eq. (10), we see that for the current parameter values in TABLE 1, $\phi$ takes non-negligible value only when $\sigma_{j}z_{j}\lesssim 1$ for each $j$. It is necessary to ensure that the grid points cover the whole of such a region with a sufficiently fine interval. Currently, as $z_{j}$ ranges from $-N\eta$ to $+N\eta$ with step size $\eta_{j}$, $\eta_{j}$ and $\sigma_{j}$ must satisfy $\sigma_{j}\eta_{j}\ll 1$ and $\sigma_{j}\times N\eta_{j}\gg 1$. Therefore, in general, we need the adjustment that $\eta_{j}$ increases when $\sigma$ decreases and vice versa. Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{j}$ set to some constant, and thus we did so. In the cases with $d=10,11$, $\eta_{j}$ is adjusted as
$\eta_{j}=\frac{a}{\sigma_{j}}+b$ (17)
To set the constants $a$ and $b$, we ran TT-based option pricing for several values of $\sigma_{j}$ adjusting $\eta_{j}$, found the pairs $(\sigma_{j},\eta_{j})$ that yields the accurate price, and adjust $a$ and $b$ so that Eq. (17) approximately reproduces such $\eta_{j}$ from $\sigma_{j}$.
Also when $\vec{p}=\vec{S}_{0}$, we set $\eta_{j}$ to some constant in almost all the test cases. Only for $d=11$, we try the adjustment that
$\eta_{j}=\frac{\alpha}{S_{j,0}}+\beta$ (18)
with $\alpha$ and $\beta$ chosen according to the option pricing result for several $\eta_{j}$ and $\sigma_{j}$. In general, larger $S_{j,0}$, which means larger $x_{j,0}$, makes $\phi$ as a function of $z_{j}$ oscillate more rapidly, and thus we need finer grid points to resolve such an oscillation.
Note that when $\eta_{j}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$.

### v3

@@ I Introduction
In light of this, applications of quantum-inspired algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present or the near future [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9] learning to the Fourier transform (FT)-based option pricing method [10, 11], following the original proposal in [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [12, 13] as reviews. and recently, it has also been utilized in various fields such as machine learning [14, 15], quantum field theory [16, 17], and differential equations [7, 18]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [19, 20, 17, 21] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
To address these issues, we focus on another usage of tensor networks that can embed parameter dependence [22, 23, 24] to make FT-based option pricing more efficient. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We use these tensor networks to perform fast option pricing in response to various parameter changes. To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice. Indeed, the computational complexity of our proposed method is better than that of Monte Carlo by a factor of $O(10^{4})$ for $d<10$ and by a factor of $O(10^{3})$ for about $d=10$ or $11$ (refer to FIG. 2). We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
@@ III.1 Fourier transform-based option pricing
where $\Sigma:=(\sigma_{m}\sigma_{n}\rho_{mn}T)_{mn}$ is the covariance matrix of $\vec{x}$ and $\vec{\mu}:=\vec{x}_{0}+\left(rT-\frac{1}{2}\sigma_{1}^{2}T,\cdots,rT-\frac{1}{2}\sigma_{d}^{2}T\right)$. Note that, in Eq. (6), we denote the option price by $V(\vec{p})$, indicating its dependence on the parameter $\vec{p}$ such as the volatilities $\vec{\sigma}=(\sigma_{1},\cdots,\sigma_{d})$ and the present asset prices $\vec{S}_{0}$.
@@
$\phi(\vec{z})=\exp\left(i\sum_{m=1}^{d}z_{m}\mu_{m}-\frac{T}{2}\sum_{m=1}^{d}\sum_{k=1}^{d}\sigma_{m}\sigma_{k}z_{m}z_{k}\rho_{mk}\right).$ (10)
@@
$\hat{v}_{\min}(\vec{z})=-\frac{K^{1+i\sum_{m=1}^{d}z_{m}}}{(-1)^{d}\left(1+i\sum_{m=1}^{d}z_{m}\right)\prod_{m=1}^{d}(iz_{m})}.$ (12)
@@
where $\eta_{m}$ is the integration step size in the $m$-th direction. $\eta_{m}$ is a hyperparameter that must be appropriately determined according to the parameter $\vec{p}$ and thus written as $\eta_{m}(\vec{p})$ in (14). In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec. V for the detail. $\Delta_{\rm vol}:=\Pi_{m=1}^{d}\eta_{m}$ is the volume element.
@@ III.2 Fourier transform-based option pricing with tensor trains
$V(\vec{\sigma},\vec{S}_{0})\simeq\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\tilde{\phi}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}}\Delta_{\mathrm{vol}}.$ (15)
Thanks to TCI, we can obtain the approximate TTs avoiding the evaluations of $\phi$ and $\hat{v}$ at all the grid points. Besides, given the TTs, we can compute the sum in (15) as the contraction of two TTs without exponentially many iterations: with the bond dimensions at most $\chi$, the number of multiplications and additions is of order $O(dN\chi^{3})$.
@@ III.3 Monte Carlo-based option pricing
IV Learning parameter dependence with tensor networks
@@ IV.1 Outline
Option prices $V(\vec{p})$ depend on input parameters $\vec{p}$ such as the volatilities $\vec{\sigma}$ and the present asset prices $\vec{S}_{0}$. In the rapidly changing financial market, these input parameters vary from time to time, which causes the change of the option price. Therefore, if we have a function, i.e., tensor networks, that efficiently outputs an accurate approximation of the option price for various values of the input parameters, it provides a large benefit to practical business.
Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set. Using TCI, we obtain the TTs to approximate $\phi$ and $\hat{v}$, incorporating the parameter dependence of these functions. The outline is illustrated in Fig. 1. Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$. Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$. Then, as illustrated in Fig. 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$. We reduce the bond dimensions of these TTs using SVD. We then contract adjacent core tensors pairwise and obtain two matrix product operators (MPOs) $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$. We contract these MPOs along the index vector $\vec{j}$ as
$K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (17)
and optimize the bond dimensions of this resultant MPO $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ using SVD. Having this MPO, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig. 1 (b). By fixing the index $\vec{k}$ of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
$V(\vec{p}_{\mathrm{gr},\vec{\hat{k}}})\simeq\frac{e^{-rT}}{2\pi}K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}\Delta_{\mathrm{vol}}.$ (18)
@@
Figure 1: Fast option pricing based on TNs proposed in this study. In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD. We then contract the tensors associated with $p_{m}$ and $z_{m}$, resulting in two MPOs. We then take the contraction of these two MPOs with respect to the index vector $\vec{j}$ and compress the bond dimension of the resulting MPO using SVD. In (b), we use this MPO to perform fast option pricing for a specified parameter $\vec{p}$.
@@ IV.2 Computational complexity
The computational complexity of TT-based option pricing, involving computation for obtaining the specific tensor components $K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ is $O(d\chi_{K}^{2})$ [25]. Here, we denote the maximum bond dimensions as $\chi_{K}$ In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations. Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec. V.
Here, we ignore the computational complexity of all the processes in FIG. 1 (a) and consider that in FIG. 1 (b) after we get the MPO only. This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing. As discussed in Sec. VI, we can reasonably find such a situation in practice.
@@ V.1 Details
Ranges of ${\sigma_{m}}$ and ${S_{m,0}}$
With respect to $\vec{p}$, on which the TTs learn the dependence of the functions $\phi$ and $\hat{v}$, we take the two test cases: $\vec{p}$ is the volatilities $\vec{\sigma}$ or the present asset prices $\vec{S}_{0}$. In the proposed method, we need to set the range in the space of $\vec{p}$ and the grid points in it. For each volatility $\sigma_{m}$, we set the range to $\sigma_{m}\in[0.15,0.25]$, where the center $\sigma_{m}=0.2$ is a typical value of Nikkei Stock Average Volatility Index [28] and the range width ±$\pm 0.05$ covers the changes in this index on most days. For each $S_{m,0}$, we set the range to $S_{m,0}\in[90,120]$, which corresponds to the 20% variation of the asset price centered at 100. The lower bound is set to not 80 but 90 because the price of the option we take as an example is negligibly small for $S_{0}<90$. For both $\sigma_{m}$ and $S_{m,0}$, we set 100 equally spaced grid points in the range, and so the total number of the grid points in the space of $\vec{p}$ is $100^{d}$.
@@
The spread of the characteristic function changes depending on the values of the parameters $\vec{p}$ and the number of assets. Accordingly, it is necessary to adjust the grid interval $\eta_{m}$ for integration in the wavenumber space.
For example, consider the case where $\vec{p}=\vec{\sigma}$. From Eq. (10), we see that for the current parameter values in TABLE 1, $\phi$ takes non-negligible value only when $\sigma_{m}z_{m}\lesssim 1$ for each $m$. It is necessary to ensure that the grid points cover the whole of such a region with a sufficiently fine interval. Currently, as $z_{m}$ ranges from $-N\eta$ to $+N\eta$ with step size $\eta_{m}$, $\eta_{m}$ and $\sigma_{m}$ must satisfy $\sigma_{m}\eta_{m}\ll 1$ and $\sigma_{m}\times N\eta_{m}\gg 1$. Therefore, in general, we need the adjustment that $\eta_{m}$ increases when $\sigma$ decreases and vice versa. Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{m}$ set to some constant, and thus we did so. In the cases with $d=10,11$, $\eta_{m}$ is adjusted as
$\eta_{m}=\frac{a}{\sigma_{m}}+b$ (19)
To set the constants $a$ and $b$, we ran TT-based option pricing for several values of $\sigma_{m}$ adjusting $\eta_{m}$, found the pairs $(\sigma_{m},\eta_{m})$ that yields the accurate price, and adjust $a$ and $b$ so that Eq. (19) approximately reproduces such $\eta_{m}$ from $\sigma_{m}$.
Also when $\vec{p}=\vec{S}_{0}$, we set $\eta_{m}$ to some constant in almost all the test cases. Only for $d=11$, we try the adjustment that
$\eta_{m}=\frac{\alpha}{S_{m,0}}+\beta$ (20)
with $\alpha$ and $\beta$ chosen according to the option pricing result for several $\eta_{m}$ and $\sigma_{m}$. In general, larger $S_{m,0}$, which means larger $x_{m,0}$, makes $\phi$ as a function of $z_{m}$ oscillate more rapidly, and thus we need finer grid points to resolve such an oscillation.
Note that when $\eta_{m}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$.